OpenStudy (princeevee):

can you hlp me Which number is between 1/6 and 0.25? A. 7/24 B. 1/7 C. 3/16 D. 2/7 (you dont have to give me the answer completely if you want)

OpenStudy (anonymous):

first you must turn 0.25 into a fraction

OpenStudy (anonymous):

if you find the lcd (lowest common denominator) you can eliminate options there

OpenStudy (anonymous):

the lcd of all these fractions is 11088

OpenStudy (princeevee):

k

OpenStudy (anonymous):

sorry, the lcd is not 11088, i had the wrong number in my head

OpenStudy (princeevee):

ok

OpenStudy (anonymous):

just use this website, it's smarter than me XD

OpenStudy (princeevee):

well you can still hlp

OpenStudy (princeevee):

ok so the lcd is *336*

OpenStudy (princeevee):

so what should i do next?

OpenStudy (anonymous):

just look at the fractions and see which ones fit in between 1/6 and 1/4

OpenStudy (princeevee):

1/7?

OpenStudy (anonymous):

take not that 1/6 = 56/336 1/4 = 84/336

OpenStudy (anonymous):

what is 1/7 when the denominator is 336?

OpenStudy (princeevee):

... i cant compute that well : /

OpenStudy (anonymous):

7/24 = 98/336 1/7 = 48/336 3/16 = 63/336 2/7 = 96/336

OpenStudy (anonymous):

which of those choices fit in between: 56/336 and 84/336 ?

OpenStudy (princeevee):

3/16

OpenStudy (anonymous):

correct

OpenStudy (princeevee):

: D

OpenStudy (princeevee):

thx